OpenStudy (anonymous):

find the supplement and complement of θ=3π/5

OpenStudy (lgbasallote):

supplement would be \(\pi - 3\pi / 5\) complement will be \(\pi/2 - 3\pi/5\)

OpenStudy (anonymous):

Supplement means two angles added up to get 180 degrees, Complement means two angles added up to get 90 degrees: The question wants the angles in radians, so if you remember from your trig class: \[90^o = \pi/2\] \[180^o = \pi\] So to find the supplementary angle to (3pi/5), we set up the algebraic equation: (theta) + 3pi/5 = pi (theta) = 2pi/5 So the supplementary angle is 2pi/5. Next is the complimentary angle, we set up the equation: (theta) + 3pi/5 = pi/2 (theta) = -pi/10 So the complementary angle is -pi/10